# Predictive Communications for Low-Altitude Networks

### v2

@@ I Introduction
The emergence of the low-altitude economy, an ecosystem of services conducted by unmanned aerial vehicles (UAVs) and electric vertical takeoff and landing (eVTOL) aircraft, is set to redefine urban and rural infrastructure [1, 2]. From automated logistics to real-time infrastructure inspection, these services depend on the seamless operation of extensive low-altitude networks. As these networks form a dynamic aerial layer above the terrestrial systems [3, 4], a communication paradigm that ensures efficient, reliable, and secure operation becomes paramount.
@@
However, low-altitude networks present a unique hybrid challenge that is not addressed by existing paradigms. They are far more dynamic than conventional terrestrial networks [5, 6], with high-speed, 3D [^1] Throughout this paper, we use the notation xD to denote an x-dimensional space. mobility that induces drastic channel variations. However, they are more constrained and operate in a much more complex environment than satellite networks. While space systems [7, 8] benefit from near-perfect trajectory predictability in an electromagnetically sparse void, low-altitude aircraft face unpredictable wind, dynamic obstacles, and a dense, interference-coupled airspace that invalidates the assumptions of traditional satellite communication design.
The opportunity to master this environment lies in harnessing a powerful, yet underutilized, resource: imperfect but statistically robust predictability. This foresight is derived from two pillars: the a priori knowledge of mission-driven flight plans filed with a central authority [9, 10, 11], as illustrated in Fig. 1, and the quasi-static nature of the large-scale radio environment, which can be captured in models like radio maps [12, 13, 14], as illustrated in Fig. 2. This available foresight allows us to shift from a reactive to a proactive design philosophy fundamentally.
@@ II-B Pillar 2: Radio Environment Predictability
  • Advanced radio-environment models now provide fine-grained, city-scale accuracy. State-of-the-art radio maps generated via ray tracing, measurement-driven interpolation, or high-fidelity digital twins can deliver metre-level resolution across the 3D workspace [12, 13, 14].
@@ V Case Study and Security Extensions
This section validates our layered predictive framework through a case study on interference mitigation [15], with key results illustrated in Fig. 5. First, the large-scale layer demonstrates strategic adaptability, exploiting aircraft mobility for delay-tolerant tasks (Fig. 5(a.2)) while using multi-hop relays for delay-sensitive tasks (Fig. 5(a.3)). The middle-scale layer then performs tactical timing coordination based on channel dynamics, scheduling a high-power, rapid transmission for the degrading first two hops in Fig. 5(a.2) while using a low-power, longer transmission for the improving last hop. The small-scale layer ensures fine-grained interference control, strategically reducing transmit power when aircraft operate near sensitive ground nodes. Compared with the classical aggregate-based baseline (routing on a static graph with locally coordinated handovers) and the space-time-based baseline (routing on a space-time graph with fixed, equally spaced handover times), the proposed layered strategy consistently achieves a larger reduction in aggregate interference power, as validated in Fig. 5(b), with this advantage becoming more pronounced as network load increases. Note that goal-oriented network formulations reduce the computational burden and keep centralized implementation tractable.

### v3

@@ I Introduction
The emergence of the low-altitude economy, an ecosystem of services conducted by unmanned aerial vehicles (UAVs) and electric vertical takeoff and landing (eVTOL) aircraft, is set to redefine urban and rural infrastructure [SonLinWanSun:M25, JiaLiZhuLi:M25]. From automated logistics to real-time infrastructure inspection, these services depend on the seamless operation of extensive low-altitude networks. As these networks form a dynamic aerial layer above the terrestrial systems [VaeLinZha:J24, SonZenYanRen:J25], a communication paradigm that ensures efficient, reliable, and secure operation becomes paramount.
@@
However, low-altitude networks present a unique hybrid challenge that is not addressed by existing paradigms. They are far more dynamic than conventional terrestrial networks [SinKumMajSat:J25, HanJiaYu:J25], with high-speed, 3D [^1] Throughout this paper, we use the notation xD to denote an x-dimensional space. mobility that induces drastic channel variations. However, they are more constrained and operate in a much more complex environment than satellite networks. While space systems [HanXuZhaWan:J23, NtoTomRui:C24] benefit from near-perfect trajectory predictability in an electromagnetically sparse void, low-altitude aircraft face unpredictable wind, dynamic obstacles, and a dense, interference-coupled airspace that invalidates the assumptions of traditional satellite communication design.
The opportunity to master this environment lies in harnessing a powerful, yet underutilized, resource: imperfect but statistically robust predictability. This foresight is derived from two pillars: the a priori knowledge of mission-driven flight plans filed with a central authority [LiChe:J24b, LiuXuWanChe:J20, LiChe:J24], as illustrated in Fig. 1, and the quasi-static nature of the large-scale radio environment, which can be captured in models like radio maps [HeAiGuaWan:J19, SunChe:J24, WanZhaNieYu:M25], as illustrated in Fig. 2. This available foresight allows us to shift from a reactive to a proactive design philosophy fundamentally.
@@ II-B Pillar 2: Radio Environment Predictability
  • Advanced radio-environment models now provide fine-grained, city-scale accuracy. State-of-the-art radio maps generated via ray tracing, measurement-driven interpolation, or high-fidelity digital twins can deliver metre-level resolution across the 3D workspace [HeAiGuaWan:J19, SunChe:J24, WanZhaNieYu:M25].
@@ V Case Study and Security Extensions
This section validates our layered predictive framework through a case study on interference mitigation [LiChe:J26], with key results illustrated in Fig. 5. First, the large-scale layer demonstrates strategic adaptability, exploiting aircraft mobility for delay-tolerant tasks (Fig. 5(a.2)) while using multi-hop relays for delay-sensitive tasks (Fig. 5(a.3)). The middle-scale layer then performs tactical timing coordination based on channel dynamics, scheduling a high-power, rapid transmission for the degrading first two hops in Fig. 5(a.2) while using a low-power, longer transmission for the improving last hop. The small-scale layer ensures fine-grained interference control, strategically reducing transmit power when aircraft operate near sensitive ground nodes. Compared with the classical aggregate-based baseline (routing on a static graph with locally coordinated handovers) and the space-time-based baseline (routing on a space-time graph with fixed, equally spaced handover times), the proposed layered strategy consistently achieves a larger reduction in aggregate interference power, as validated in Fig. 5(b), with this advantage becoming more pronounced as network load increases. Note that goal-oriented network formulations reduce the computational burden and keep centralized implementation tractable.
@@ VII Conclusion
Biographies
Junting Chen (S’11–M’16) received the Ph.D. degree in Electronic and Computer Engineering from the Hong Kong University of Science and Technology (HKUST), Hong Kong SAR China, in 2015, and the B.Sc. degree in Electronic Engineering from Nanjing University, Nanjing, China, in 2009. From 2014–2015, he was a visiting student with the Wireless Information and Network Sciences Laboratory at MIT, Cambridge, MA, USA.
He is an Assistant Professor with the School of Science and Engineering and the Shenzhen Future Network of Intelligence Institute (FNii–Shenzhen) at The Chinese University of Hong Kong, Shenzhen (CUHK–Shenzhen), Guangdong, China. Prior to joining CUHK–Shenzhen, he was a Postdoctoral Research Associate with the Ming Hsieh Department of Electrical Engineering, University of Southern California (USC), Los Angeles, CA, USA, from 2016–2018, and with the Communication Systems Department of EURECOM, Sophia–Antipolis, France, from 2015–2016. His research interests include channel estimation, MIMO beamforming, machine learning, and optimization for wireless communications and localization. His current research focuses on radio map sensing, construction, and application for wireless communications. Dr. Chen was a recipient of the HKTIIT Post-Graduate Excellence Scholarships in 2012. He was nominated as the Exemplary Reviewer of IEEE Wireless Communications Letters in 2018. His paper received the Charles Kao Best Paper Award from WOCC 2022.
Bowen Li (Member, IEEE) received the Ph.D. degree in Computer and Information Engineering from The Chinese University of Hong Kong, Shenzhen (CUHK–Shenzhen), China, in 2025. He is currently a Postdoctoral Research Associate with Department of Computer and Information Science, Linköping University. His research interests primarily include predictive communications, semantic communications, and signal processing for low-altitude networks.
Hao Sun received the B.S. degree from the University of Electronic Science and Technology of China (UESTC), Chengdu, China, in 2018. He received the Ph.D. degree in Computer and Information Engineering from The Chinese University of Hong Kong, Shenzhen (CUHK–Shenzhen), Guangdong, China, in 2025. He is currently a Postdoctoral Fellow with the Department of Electrical Engineering, City University of Hong Kong, Hong Kong. His research interests include matrix completion, tensor decomposition, and generative AI with applications with applications in wireless communication and signal processing.
Shuguang Cui (Fellow, IEEE) received the Ph.D. degree from Stanford University in 2005. He is currently a X. Q. Deng Presidential Chair Professor at The Chinese University of Hong Kong, Shenzhen, China. His current research interests include data driven large-scale information analysis and system design. He was selected as the Thomson Reuters Highly Cited Researcher and listed in the Worlds’ Most Influential Scientific Minds by ScienceWatch in 2014. He was a recipient of the IEEE SP Society 2012 and ComSoc 2023 Marconi Best Paper Awards. He is a Member of Both Royal Society of Canada and Canadian Academy of Engineering.
Nikolaos Pappas (Senior Member, IEEE) is currently an Associate Professor with the Department of Computer and Information Science, Linköping University, Linköping, Sweden. His research interests include the field of wireless communication networks, with an emphasis on semantics-aware communications, energy harvesting networks, age of information, and stochastic geometry. He has served as the Symposium Co-Chair for the IEEE International Conference on Communications in 2022. He was the General Chair for the 23rd International Symposium on Modeling and Optimization in Mobile, Ad hoc, and Wireless Networks (WiOpt 2025). He is an Area Editor of IEEE OPEN JOURNAL OF THE COMMUNICATIONS SOCIETY and an Expert Editor of invited papers of IEEE COMMUNICATIONS LETTERS. He is an Associate Editor of four IEEE TRANSACTIONS and journals.